# Magnitude homology of real hyperplane arrangements

### 04 Apr 2026 v1

@@ Appendix A Computer experiments
The following examples are computed by using SageMath [28].
@@
We list the results of magnitudes of simplicial arrangements $\mathcal{A}(n,k)$ in Grünbaum’s catalogue [14] for $n\leq 13$. Note that $\mathcal{A}(n,0),\mathcal{A}(6,1),\mathcal{A}(9,1)$ are recorded in Examples 3.10 and 3.11. We use the database of [6] to pull the list of normal vectors.
@@
Consider the non-simplicial arrangement $\mathcal{A}$ defined by $xyz(x+z)(y+z)$, which is also the graphic arrangement $\mathcal{A}$ of the graph $K_{4}\setminus e$. Note that $K_{4}\setminus e$ is chordal.
@@
where the numerator $P(q)=96q^{8}-108q^{7}+48q^{6}+36q^{5}+36q^{3}+48q^{2}-108q+96$ and the denominator $Q(q)=q^{17}+3q^{16}+5q^{15}+7q^{14}+8q^{13}+9q^{12}+11q^{11}+13q^{10}+15q^{9}+15q^{8}+13q^{7}+11q^{6}+9q^{5}+8q^{4}+7q^{3}+5q^{2}+3q+1$.
@@
Consider the bracelet arrangement $\mathcal{A}$ defined by $x_{1}x_{2}x_{3}(x_{1}+x_{4})(x_{2}+x_{4})(x_{3}+x_{4})(x_{1}+x_{2}+x_{4})(x_{1}+x_{3}+x_{4})(x_{2}+x_{3}+x_{4})$ (see Example 1.4 of [7]),
@@
where the numerator $P(q)=102q^{8}-24q^{7}-48q^{6}-54q^{5}+264q^{4}-54q^{3}-48q^{2}-24q+102$ and the denominator $Q(q)=q^{17}+4q^{16}+8q^{15}+11q^{14}+12q^{13}+13q^{12}+16q^{11}+20q^{10}+23q^{9}+23q^{8}+20q^{7}+16q^{6}+13q^{5}+12q^{4}+11q^{3}+8q^{2}+4q+1$.
@@
We collect here some results of magnitude Betti numbers using SageMath codes of [16].
@@
Table 1. Magnitude Betti numbers $\beta_{k,l}$ for the Boolean arrangement $Bl_{2}=U_{2,2}$ defined by $xy=0$.
@@
Table 2. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $A_{2}=U_{2,3}$.
@@
Table 3. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $B_{2}=U_{2,4}$.
@@
Table 4. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $A_{3}$.
@@
Table 5. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $A_{4}$.
@@
Table 6. Magnitude Betti numbers $\beta_{k,l}$ for the generic arrangement $U_{3,4}$ defined by $xyz(x+y+z)$ Example A.2.
@@
Table 7. Magnitude Betti numbers $\beta_{k,l}$ for the generic arrangement $U_{4,5}$.
@@
Table 8. Magnitude Betti numbers $\beta_{k,l}$ for the graphic arrangement of $K_{4}\setminus e$ Example A.3
@@
Table 9. Magnitude Betti numbers $\beta_{k,l}$ for $\mathcal{A}(7,1)$
@@
Table 10. Magnitude Betti numbers $\beta_{k,l}$ for $\mathcal{A}(8,1)$
@@
Table 11. Magnitude Betti numbers $\beta_{k,l}$ for $\mathcal{A}(9,1)$
Reference
  • [1] T. Abe, T. Horiguchi, M. Masuda, S. Murai, and T. Sato (2020) Hessenberg varieties and hyperplane arrangements. J. Reine Angew. Math. 764, pp. 241–286. External Links: ISSN 0075-4102,1435-5345, Document, Link, MathReview (Marko Radovanović) Cited by: §2.7.
  • [2] M. Aguiar and S. Mahajan (2017) Topics in hyperplane arrangements. Mathematical Surveys and Monographs, Vol. 226, American Mathematical Society, Providence, RI. External Links: ISBN 978-1-4704-3711-4, Document, Link, MathReview (Tom C. Braden) Cited by: §2.6, Proposition 2.2, Remark 2.25, Theorem 2.33.
  • [3] A. Björner, P. H. Edelman, and G. M. Ziegler (1990) Hyperplane arrangements with a lattice of regions. Discrete Comput. Geom. 5 (3), pp. 263–288. External Links: ISSN 0179-5376,1432-0444, Document, Link, MathReview (Hiroaki Terao) Cited by: §2.7, Proposition 2.36.
  • [4] A. Björner, M. Las Vergnas, B. Sturmfels, N. White, and G. M. Ziegler (1999) Oriented matroids. Second edition, Encyclopedia of Mathematics and its Applications, Vol. 46, Cambridge University Press, Cambridge. External Links: ISBN 0-521-77750-X, Document, Link, MathReview Entry Cited by: Theorem 1.1, Proposition 2.30, Proposition 2.34, Corollary 2.35.
  • [5] S. Brenner, J. Cardinal, T. McConville, A. Merino, and T. Mütze (2026) Combinatorial generation via permutation languages. VII. Supersolvable hyperplane arrangements. European J. Combin. 135, pp. Paper No. 104367. External Links: ISSN 0195-6698,1095-9971, Document, Link, MathReview Entry Cited by: §2.5.
  • [6] M. Cuntz, S. Elia, and J. Labbé (2022) Congruence normality of simplicial hyperplane arrangements via oriented matroids. Ann. Comb. 26 (1), pp. 1–85. External Links: ISSN 0218-0006,0219-3094, Document, Link, MathReview (Hery Randriamaro) Cited by: Example A.1.
  • [7] G. Denham, G. Dorpalen-Barry, and N. Proudfoot (2026) Cleanliness and the varchenko-gelfand algebra. arXiv preprint arXiv:2512.10077v2. Cited by: Example A.5.
  • [8] G. Denham and P. Hanlon (1999) Some algebraic properties of the Schechtman-Varchenko bilinear forms. In New perspectives in algebraic combinatorics (Berkeley, CA, 1996–97), Math. Sci. Res. Inst. Publ., Vol. 38, pp. 149–176. External Links: ISBN 0-521-77087-4, MathReview (P. Orlik) Cited by: §2.6.
  • [9] R. Diestel ([2025] ©2025) Graph theory. Sixth edition, Graduate Texts in Mathematics, Vol. 173, Springer, Berlin. External Links: ISBN 978-3-662-70106-5; 978-3-662-70107-2, MathReview Entry Cited by: §2.5.
  • [10] P. H. Edelman and J. W. Walker (1985) The homotopy type of hyperplane posets. Proc. Amer. Math. Soc. 94 (2), pp. 221–225. External Links: ISSN 0002-9939,1088-6826, Document, Link, MathReview (Udo Pachner) Cited by: Proposition 2.34.
  • [11] P. H. Edelman (1984) A partial order on the regions of ${\bf R}^{n}$ dissected by hyperplanes. Trans. Amer. Math. Soc. 283 (2), pp. 617–631. External Links: ISSN 0002-9947,1088-6850, Document, Link, MathReview (Robert A. Proctor) Cited by: Corollary 2.35.
  • [12] L. Giordani, T. Möller, P. Mücksch, and G. Roehrle (2025) On connected subgraph arrangements. arXiv preprint arXiv:2502.18144v1. Cited by: §2.7.
  • [13] K. Gomi (2025) Magnitude homology of geodesic space. arXiv preprint arXiv:1902.07044v3. Cited by: Proposition 2.21.
  • [14] B. Grünbaum (2009) A catalogue of simplicial arrangements in the real projective plane. Ars Math. Contemp. 2 (1), pp. 1–25. External Links: ISSN 1855-3966,1855-3974, Document, Link, MathReview Entry Cited by: Example A.1, §1, §3.3.
  • [15] Y. Gu (2018) Graph magnitude homology via algebraic morse theory. arXiv preprint arXiv:1809.07240v1. Cited by: Example 2.16.
  • [16] R. Hepworth and S. Willerton (2017) Categorifying the magnitude of a graph. Homology Homotopy Appl. 19 (2), pp. 31–60. External Links: ISSN 1532-0073,1532-0081, Document, Link, MathReview (Haimiao Chen) Cited by: Example A.6, §1, §2.3, Proposition 2.10, Theorem 2.12, Theorem 2.12, Proposition 2.8, Proposition 2.9, §4.5.
  • [17] W. Hochstättler and V. Welker (2019) The Varchenko determinant for oriented matroids. Math. Z. 293 (3-4), pp. 1415–1430. External Links: ISSN 0025-5874,1432-1823, Document, Link, MathReview (Daryl Funk) Cited by: §2.6.
  • [18] M. Jambu and L. Paris (1995) Combinatorics of inductively factored arrangements. European J. Combin. 16 (3), pp. 267–292. External Links: ISSN 0195-6698,1095-9971, Document, Link, MathReview (Hiroaki Terao) Cited by: §2.7.
  • [19] R. Kaneta and M. Yoshinaga (2021) Magnitude homology of metric spaces and order complexes. Bull. Lond. Math. Soc. 53 (3), pp. 893–905. External Links: ISSN 0024-6093,1469-2120, Document, Link, MathReview (David Matthew Freeman) Cited by: §2.3, §2.3, Theorem 2.20, Theorem 2.23.
  • [20] V. Körber, T. Schnieders, J. Stricker, and J. Walizadeh (2025) Hamiltonian cycles in simplicial and supersolvable hyperplane arrangements. arXiv preprint arXiv:2508.14538v2. Cited by: §2.5.
  • [21] T. Leinster and M. W. Meckes (2017) The magnitude of a metric space: from category theory to geometric measure theory. In Measure theory in non-smooth spaces, Partial Differ. Equ. Meas. Theory, pp. 156–193. External Links: ISBN 978-3-11-055082-5; 978-3-11-055083-2, MathReview (Tobias W. Fritz) Cited by: §1.
  • [22] T. Leinster and M. Meckes (2023) Spaces of extremal magnitude. Proc. Amer. Math. Soc. 151 (9), pp. 3967–3973. External Links: ISSN 0002-9939,1088-6826, Document, Link, MathReview (Shlomo Reisner) Cited by: §3.1.
  • [23] T. Leinster and M. Shulman (2021) Magnitude homology of enriched categories and metric spaces. Algebr. Geom. Topol. 21 (5), pp. 2175–2221. External Links: ISSN 1472-2747,1472-2739, Document, Link, MathReview (Ahmet A. Husainov) Cited by: §1.
  • [24] T. Leinster (2013) The magnitude of metric spaces. Doc. Math. 18, pp. 857–905. External Links: ISSN 1431-0635,1431-0643, MathReview (J. Böhm) Cited by: §1.
  • [25] T. Leinster (2019) The magnitude of a graph. Math. Proc. Cambridge Philos. Soc. 166 (2), pp. 247–264. External Links: ISSN 0305-0041,1469-8064, Document, Link, MathReview (Lorenzo Traldi) Cited by: §1, §2.2, Proposition 2.3, Proposition 2.5, Theorem 2.6, Proposition 2.7.
  • [26] S. Oh, A. Postnikov, and H. Yoo (2008) Bruhat order, smooth Schubert varieties, and hyperplane arrangements. J. Combin. Theory Ser. A 115 (7), pp. 1156–1166. External Links: ISSN 0097-3165,1096-0899, Document, Link, MathReview (David E. Anderson) Cited by: §2.7.
  • [27] P. Orlik and H. Terao (1992) Arrangements of hyperplanes. Grundlehren der mathematischen Wissenschaften [Fundamental Principles of Mathematical Sciences], Vol. 300, Springer-Verlag, Berlin. External Links: ISBN 3-540-55259-6, Document, Link, MathReview (Michel Yves Jambu) Cited by: §2.4.
  • [28] The Sage Developers (2025) Sagemath, the Sage Mathematics Software System (Version 10.8). Note: https://www.sagemath.org Cited by: Appendix A.
  • [29] L. Solomon (1966) The orders of the finite Chevalley groups. J. Algebra 3, pp. 376–393. External Links: ISSN 0021-8693, Document, Link, MathReview (T. Ono) Cited by: Proposition 2.37.
  • [30] R. P. Stanley (2007) An introduction to hyperplane arrangements. In Geometric combinatorics, IAS/Park City Math. Ser., Vol. 13, pp. 389–496. External Links: ISBN 978-0-8218-3736-8; 0-8218-3736-2, Document, Link, MathReview Entry Cited by: §2.4, Remark 2.25.
  • [31] A. Varchenko (1993) Bilinear form of real configuration of hyperplanes. Adv. Math. 97 (1), pp. 110–144. External Links: ISSN 0001-8708,1090-2082, Document, Link, MathReview (Hiroaki Terao) Cited by: §2.6, Theorem 2.33.
  • [32] Y. Yagi and M. Yoshinaga (2026) Reconstrucion of oriented matroids from varchenko-gelfand algebras. arXiv preprint arXiv:2509.19905v3. Cited by: §1.
  • [33] T. Zaslavsky (1975) Facing up to arrangements: face-count formulas for partitions of space by hyperplanes. Mem. Amer. Math. Soc. 1, pp. vii+102. External Links: ISSN 0065-9266,1947-6221, Document, Link, MathReview (E. Jucovič) Cited by: §2.4, Remark 2.25.

### 01 May 2026 v3

@@ Appendix A Computer experiments
The following examples are computed with SageMath [29].
@@
We list the magnitudes of simplicial arrangements $\mathcal{A}(n,k)$ in Grünbaum’s catalogue [15] for $n\leq 13$. Note that $\mathcal{A}(n,0),\mathcal{A}(6,1),\mathcal{A}(9,1)$ are recorded in Examples 3.12 and 3.13. We use the database of [6] to obtain the list of normal vectors.
@@
Consider the non-simplicial arrangement $\mathcal{A}$ defined by $xyz(x+z)(y+z)$, which is also the graphic arrangement of the graph $K_{4}\setminus e$. Note that $K_{4}\setminus e$ is chordal.
@@
where the numerator is $P(q)=96q^{8}-108q^{7}+48q^{6}+36q^{5}+36q^{3}+48q^{2}-108q+96$ and the denominator is $Q(q)=q^{17}+3q^{16}+5q^{15}+7q^{14}+8q^{13}+9q^{12}+11q^{11}+13q^{10}+15q^{9}+15q^{8}+13q^{7}+11q^{6}+9q^{5}+8q^{4}+7q^{3}+5q^{2}+3q+1$.
@@
Consider the bracelet arrangement $\mathcal{A}$ defined by $x_{1}x_{2}x_{3}(x_{1}+x_{4})(x_{2}+x_{4})(x_{3}+x_{4})(x_{1}+x_{2}+x_{4})(x_{1}+x_{3}+x_{4})(x_{2}+x_{3}+x_{4})$ (see Example 1.4 of [7]).
@@
where the numerator is $P(q)=102q^{8}-24q^{7}-48q^{6}-54q^{5}+264q^{4}-54q^{3}-48q^{2}-24q+102$ and the denominator is $Q(q)=q^{17}+4q^{16}+8q^{15}+11q^{14}+12q^{13}+13q^{12}+16q^{11}+20q^{10}+23q^{9}+23q^{8}+20q^{7}+16q^{6}+13q^{5}+12q^{4}+11q^{3}+8q^{2}+4q+1$.
@@
We collect some results here on magnitude Betti numbers computed with the SageMath code of [17].
@@
Table 1. Magnitude Betti numbers $\beta_{k,\ell}$ for the Boolean arrangement $Bl_{2}=U_{2,2}$ defined by $xy=0$.
@@
Table 2. Magnitude Betti numbers $\beta_{k,\ell}$ for the Coxeter arrangement $A_{2}=U_{2,3}$.
@@
Table 3. Magnitude Betti numbers $\beta_{k,\ell}$ for the Coxeter arrangement $B_{2}=U_{2,4}$.
@@
Table 4. Magnitude Betti numbers $\beta_{k,\ell}$ for the Coxeter arrangement $A_{3}$.
@@
Table 5. Magnitude Betti numbers $\beta_{k,\ell}$ for the Coxeter arrangement $A_{4}$.
@@
Table 6. Magnitude Betti numbers $\beta_{k,\ell}$ for the generic arrangement $U_{3,4}$ defined by $xyz(x+y+z)$; see Example A.2.
@@
Table 7. Magnitude Betti numbers $\beta_{k,\ell}$ for the generic arrangement $U_{4,5}$.
@@
Table 8. Magnitude Betti numbers $\beta_{k,\ell}$ for the graphic arrangement of $K_{4}\setminus e$; see Example A.3.
@@
Table 9. Magnitude Betti numbers $\beta_{k,\ell}$ for $\mathcal{A}(7,1)$
@@
Table 10. Magnitude Betti numbers $\beta_{k,\ell}$ for $\mathcal{A}(8,1)$
@@
Table 11. Magnitude Betti numbers $\beta_{k,\ell}$ for $\mathcal{A}(9,1)$